Given complex numbers
u = 3/2 - 2i and v = 5/3 + 3i
Compute the combination u + v

Answer by richwmiller(17219)

u = 3/2 - 2i and v = 5/3 + 3i
u +v= 3/2 - 2i + 5/3 + 3i
u+v = 9/6 - 2i +10/6 + 3i
19/6+i
3 1/6+i